OpenStudy (anonymous):

solve the equation. sqrt(p+9) +3=p

OpenStudy (zehanz):

subtract 3 from both sides: \[\sqrt{p+9}=p-3\]Squaring both sides will get you to the anser, but beware: squaring can also introduce false solutions. To avoid these, do this: 1. p + 1 >= 0, due to the root. 2. p - 3 >= 0, because it is the outcome of the root. Combine these to restrictions to see which values of p are valid. Now go on squaring and solving. Should this yield some value of p that doesn't satisfy conditions 1 and 2, then this is not a valid solution.

OpenStudy (anonymous):

i am lost on how to square them.

OpenStudy (zehanz):

Squaring the root cancels it:\[p+9=(p-3)^2\]Now expand the right side.